A total of $700 is invested for one year. Some of it is invested at 4% and the rest of it is
invested at 5%. At the end of the year, the total amount of interest earned is $30.
An amount of was invested at 4% and was invested at 5%

Found 2 solutions by Boreal, fractalier:
Answer by Boreal(15235)

x is invested at 4%
700-x is invested at 5%
0.04x+(700-x)(0.05)=0.04x + 35-0.05x=30
-0.01x=-5
x=500
700-x=200
$500 @ 4%=$20
200 @ 5%=$10
$30 total, check.

Answer by fractalier(6550)

Call the amount invested at 4%, x. Thus the amount invested at 5% is 700-x.
Then the setup looks like this:
(.04)x + (.05)(700-x) = 30
Multiply by 100
4x + 5(700-x) = 3000
4x + 3500 - 5x = 3000
-x = -500 and
x = $500 invested at 4%
Thus $200 is invested at 5%.